sinx+siny+sinz=0


then what is sin3x+sin3y+sin3z=

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

sin3A = 3sinA - 4sin3A, x3+y3= (x+y)(x2-xy+y2)

sin3x +sin3y +sin3z= 3(sinx+siny+sinz)-4(sin3x+sin3y+sin3z) = -4((sinx+siny)3 -3sinxsiny(sinx+siny)+sin3z)

=-4((sinx+siny+sinz)((sinx+siny)2 +sin2z -sinz(sinx+siny)) +12sinxsiny(sinx+siny)=-12sinxsinysinz

'Win exciting gifts by answering the questions on Discussion Forum.'